OpenStudy (anonymous):

This grade is really important, i would love it if someone could please help!!!! the number of buckets of paint n needed to paint a fence varies directly with the total area a of the fence and inversely with the amount of paint p in a bucket. it takes three buckets of paint to paint 72 square feet of fence. how many one gallon buckets will be needed to paint 90 square feet of fence?

jimthompson5910 (jim_thompson5910):

the number of buckets of paint n needed to paint a fence varies directly with the total area a of the fence and inversely with the amount of paint p in a bucket translates to n = k*a/p

jimthompson5910 (jim_thompson5910):

it takes three buckets of paint to paint 72 square feet of fence so we know that n = 3 and a = 72

jimthompson5910 (jim_thompson5910):

plug this in and solve for k n = k*a/p 3 = k*72/p 3p = 72k 72k = 3p k = 3p/72 k = p/24

jimthompson5910 (jim_thompson5910):

it further asks how many one gallon buckets will be needed to paint 90 square feet of fence? so we now can say that n = unknown, a = 90, p = 1 ----------------------- n = k*a/p n = (p/24)*(a/p) n = (1/24)*(90/1) n = 90/24 n = 15/4 n = 3.75 You'll need to round up because you can't use 0.75 of a bucket. So you round up to get 4

OpenStudy (anonymous):

Thank you so much!!!!

jimthompson5910 (jim_thompson5910):

yw